# EdgeAlpha: Distributed Process Discovery at

### v3

@@ IV-D Phase 1 – Event Ordering & Partial Footprint Matrices
Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Therefore, after the node detects the event, it queries for the predecessor event, for which it does not get any responses. Thus, the node sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response, including event 1. It increments the value in the corresponding cell of its local FM.
@@
When an event $e$ on node $i$ is triggered, node $i$ executes the following steps; see Fig. 2 for an example. First, node $i$ stores event $e$ locally, including its case ID, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event. When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp that is still smaller than the own one, as predecessor. We denote the node with the predecessor event as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event.
@@
If a node has to choose its predecessor from several events, it indicates communication delays which may cause incorrectly set predecessors. However, this is automatically rectified during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor. The node sets the correct successor and informs the node corresponding to the incorrect successor. This one starts the predecessor search again. The node rectifies the stored events once the correct event is identified and updates the counts in its partial FM accordingly. For this reason nodes store a partial FM that not only encodes whether a direct succession has occurred but also counts the occurrences. The rectification process works due to a single start event and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests. As soon as a node has more than one successor in a single trace the incorrectly set one is informed and an total order of events is established.
@@
Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
EdgeAlpha’s design further incorporates a sliding-window-principle for discarding old data in order to respond to the requirements of the resource-restricted sensor nodes. With this principle, as soon as a certain memory size has been reached or a certain time has been exceeded, the node discards the oldest data when adding new data. However, the sliding window approach only gets beneficial with a vast amount of data. The event logs we consider in the evaluation do not fill enough memory to utilize the sliding window.
IV-E Phase 2 – Requesting Process Data
This phase begins when an outside entity, for example, a central node, requests a process model. During this phase, we request the local FMs of all nodes and start or end events within any case, see Fig. 3. Subsequently, each node responds with its local FM. Before sending, the node changes all entries greater than 0 to 1, aggregating the data even more. This converts the local partial FM back into the conventional FM structure. Each node checks if events without a successor were recorded to determine whether its respective activity is an end activity. The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID. Further, we merge the start and end activities into two sets $T_{I}$ and $T_{O}$, following the conventions of the Alpha Miner. Next, the requesting unit proceeds with the remainder of the Alpha Miner algorithm, starting after the calculation of the FM, here $FM^{all}$.
@@ IV-E Phase 2 – Requesting Process Data
Figure 3: Phase 2 – Requesting Process Data: We request partial FMs and start/end activity flags from all nodes. Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the process model.
IV-F Most-Frequent-Predecessors (MFPs)
In Phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, we search for the predecessor by querying nodes in order of their likelihood of being the predecessor. Thus, when searching for an event’s predecessor, a node first contacts the most likely predecessor, then the second most likely one, and so on. For this, during runtime each node continuously records how often each predecessor occurs. As we show later in a dataset analysis, most activities in real-world scenarios are preceded by a typically stable and small subset of activities, this approach prioritizes querying the most frequently occurring predecessors. However, when detecting a start event, the corresponding node still has to query every node in the network to ensure it is the start event. In our evaluation in Section V we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures scalability.
IV-G Batching
@@
To further minimize communication overhead in EdgeAlpha, we introduce batching of querying predecessor events. Rather than sending individual requests for each event, nodes accumulate a number of events before dispatching them in a single message. The nodes continue to use MFP Requesting. As soon as the node finds the predecessor of an event, it removes the event from the batch. The node ceases to query additional nodes once the batch is empty.
IV-H Correctness
After introducing the algorithmic design of EdgeAlpha, we next prove its correctness. For this, we show in the following that EdgeAlpha outputs the same process models as the traditional centralized Alpha Miner algorithm.
We begin by introducing notations before defining our claim. Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes all activities in $L$. Further, let $FM^{i}$ for $i\in\{0,\dots,n-1\}$ be the partial FM computed by node $i$, and $L_{i}$ the partial event log that node $i$ has knowledge of. Since $i$ only keeps track of its predecessor events, its footprint matrix $FM^{i}$ contains only one column. In the following correctness claim, we omit start and end activities. The design of EdgeAlpha ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace. By definition, uniting all these activities into two distinct sets adheres to the Alpha Miner algorithm.
Claim 1
Proof 1
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity ID. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus proving the claim. $\Box$
IV-I Design Discussion
A major advantage of EdgeAlpha is that event logs are no longer stored in one central place. Instead each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity. In addition, we use batching, i.e. a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
Further, the design of EdgeAlpha directly applies to process mining algorithms beyond the Alpha Miner. Algorithms that start by building an FM or keeping track of direct successions in the event log can be projected onto our distribution approach. A well-known example of another algorithm is the Heuristics Miner [14]. The vast majority of results that we show in our evaluation in Section V, such as the efficiency of Phase 1 of EdgeAlpha, are not specific to the Alpha Miner and directly apply, for example, to the Heuristics Miner.
V Evaluation
@@
In this section, we evaluate EdgeAlpha in terms of correctness, accuracy, and performance. We present our implementation in Section V-A and suitable baselines in Section V-B. Section V-C introduces on which well-known datasets we conduct the experiments. We analyze these datasets in Section V-D. In Section V-E we discuss correctness experiments and the compliance between the FMs of EdgeAlpha and the traditional Alpha Miner. We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section V-F. Section V-G estimates FM sizes during an algorithm run and Section V-H discusses the overall scalability of EdgeAlpha.
V-A Implementation
@@
We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha.
V-B Baselines
@@
In our experiments, we compare EdgeAlpha with and without MFP Requesting. ”Query All” represents the upper bound where all nodes are queried, while a theoretical oracle-like approach serves as the lower bound. The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise.
EdgeAlpha’s design assumes event detection as part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning. Thus, comparing EdgeAlpha to the traditional Alpha Miner [12] or other distributed Alpha Miners such as the Map-Reduce approach [19] or the Alpha Miner in the healthcare scenario [9] is not viable. The novelty of EdgeAlpha lies in utilizing the event detection phase to pre-process events, a feature absent in other approaches.
V-C Datasets
@@
We conduct experiments on the datasets Hospital Log [13], Sepsis Cases [24] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version. [3]. These datasets naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 [25] and the Road Traffic Fine Management Process [26]. In our setting, each activity in a dataset represents a node in an IoT network.
@@
TABLE I: Dataset Characteristics (Standard deviation in parentheses if not stated otherwise).
@@
V-D Dataset Analysis
We analyze the properties of real-world event logs. Next to basic information such as the numbers of events, activities and cases and the average case length, Table I includes data that directly influences the number of requests in Phase 1. These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity. Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
@@ V-D Dataset Analysis
Even in a dataset like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event. The dataset analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events. This characteristic benefits EdgeAlpha, as self-loops do not require communication. For self-loops, too, only in case of communication delays a self-loop may be set wrongfully for which, however, EdgeAlpha provides a rectification procedure.
V-E Experimental Correctness and Fitness Analysis
@@
As a first step, we validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^3] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section V-C.
@@
Next, we examine how the merged FMs obtained by querying the nodes for their partial FMs after an increasing number of events differ from the intermediate FM produced by the traditional Alpha Miner. We compare them by using the footprint-based fitness value implemented in pm4py. Fig. 7 confirms the full compliance of EdgeAlpha’s resulting FM with the Alpha Miner’s FM across all dataset. It also demonstrates that nearly all datasets reach 90% fitness with less than 5% of events. Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, although it is considerably small, as shown in Table I. The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events there are typically present in a common event log before requesting the first process model. Moreover, our analysis suggests that reducing the sliding window for discarding old data would have minimal impact on the resulting process model, but is dependent on the current dataset.
V-F Number of Queried Nodes
To assess EdgeAlpha’s performance, we analyze its two phases separately. The event ordering phase depends highly on the number of potential predecessors a node queries, leading to significant network traffic, particularly during concurrent events, in datasets like the Hospital Log, where up to 623 nodes are queried. MFP Requesting in EdgeAlpha builds on past event data to optimize requests in the future. For the first experiments, we assume a batch size of one to focus on other characteristics. As our batch size analysis shows, the different batch sizes strongly influence the number of queried nodes.

### v4

@@ 4.2.1. Phase 1 – Event Ordering & Partial Footprint Matrices
Figure 2. Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeMiner: Without a central entity, nodes determine the order of events collaboratively using message passing. In this example, event 1 is a start event. Therefore, after the node detects the event, it queries the other nodes for the predecessor event. In our example it does not get any positive responses, and, thus, denotes it detected a start event. Upon sensing event 2, the detecting node queries for the predecessor event, receives a response, listing event 1 as predecessor. It stores this information in its local FM.
[…]
[Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeMiner ]Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeMiner: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Therefore, after the node detects the event, it queries for the predecessor event, for which it does not get any responses. Thus, the node sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response, including event 1. It increments the value in the corresponding cell of its local FM.
For simplicity, we begin by introducing a naive version of EdgeMiner, which is not communication efficient. Later, we introduce optimizations that ensure efficiency. When an event $e$ on node $i$ is triggered, node $i$ executes the following steps; see Fig. 2 for an example. First, node $i$ stores event $e$ locally, including its case ID, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event.
@@
When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp that is still smaller than the own one, as predecessor. We denote the node with the predecessor event as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event. Events which do not have successors for a long duration are flagged as end activities.
@@
If a node has to choose its predecessor from several potential predecessor events, this indicates that events are happening in quick succession and (a) communication is delayed or (b) multiple nodes are identifying predecessors in parallel. This may lead to temporarily incorrectly set predecessors. EdgeMiner automatically rectifies such inconsistencies during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor: The node sets the correct successor and informs the incorrect successor. Then, the incorrect one starts the predecessor search again. A node then corrects the counters in its local FM. The rectification process converges since each trace has a single start event, and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests.
@@
Formally, each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. In EdgeMiner, a node searches for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
4.2.2. Phase 2 – Requesting a Footprint Matrix
This phase begins when an outside entity requests the FM. During this phase, the outside entity requests the local FMs of all nodes and start and end events within any case, see Fig. 3. Subsequently, each node responds with its local FM. The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\mathbb{N}_{0}^{n\times n}$ by concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID. Next, this FM serves as a process model for conformance checking, which compares the next events, traces, or sets of traces. Further, process discovery algorithms such as the Alpha Miner (van der Aalst et al., 2004), the Heuristics Miner (Weijters et al., 2006), or the Inductive Miner (Leemans et al., 2013) can use this assembled FM for process discovery. Please note that the outside entity only receives aggregate information, i.e., the partial FMs. Raw event data is only processed on the sensor nodes, ensuring both scalability and data privacy.
@@ 4.2.2. Phase 2 – Requesting a Footprint Matrix
Figure 3. Phase 2 – Requesting a Footprint Matrix: We request partial FMs and start/end activity flags from all nodes. Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the footprint matrix.
[…]
[Phase 2 – Requesting a Footprint Matrix]We request partial FMs and start/end activity flags from all nodes. Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the footprint matrix.
4.3. Optimzaitons
Next, we introduce three optimizations: Most-Frequent-Predecessors (MFPs), batching and a sliding window to enhance efficiency.
4.3.1. Most-Frequent-Predecessors (MFPs)
In Phase 1 of EdgeMiner, a node – upon detecting an event – sends out requests to all other nodes to determine the event’s predecessor. This results in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, EdgeMiner contacts nodes in order of their likelihood of being the predecessor, i.e., it begins with the node that in the past was the most frequent predecessor according to its local FM. We show in a dataset analysis in Section 5 that activities of real-world processes statistically only have a few predecessor activities, which further only rarely change throughout the lifetime of a process. For example, even for datasets with a substantial number of activities, such as the Hospital Log (van Dongen, 2011) comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. As a result, for each event, a node can very efficiently determine the predecessor, reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events.
4.3.2. Batching
@@
To further minimize communication overhead in EdgeMiner, we batch querying of predecessor events. Rather than sending individual requests for each event, nodes accumulate a number of events before dispatching them in a single message. The nodes continue to use MFP Requesting. As soon as the node finds the predecessor of an event, it removes the event from the batch. The node ceases to query additional nodes once the batch is empty. For example, by batching predecessor queries in batches of 40 events, we show that EdgeMiner can reduce the average number of queried nodes per event to less than 2.5% of all nodes.
4.3.3. Sliding Window
To reduce resource demands, EdgeMiner utilizes a sliding window: Once the local event log reaches a given memory size or information age, the node discards it.
4.4. Correctness
After introducing the algorithmic design of EdgeMiner, we next show its correctness. We claim that merging the partial FMs in EdgeMiner outputs the same FM as when calculating the FM based on a full event log. Here, we discuss the intuition of the proof, a formal proof is in Appendix A.
By definition of the FM, there is one column and row per activity. An entry of the FM describes how often the activity of the row was followed by the activity of the column. A column of an FM therefore contains all predecessor relations of a specific activity. In EdgeMiner, the calculation of these columns is distributed to the nodes, and they are updated whenever the predecessor of an activity is found. Every predecessor relation of a node, and therefore activity, is found by the design of EdgeMiner and no direct successions are added to the FMs of multiple nodes but exactly one. Thus, locally created columns in concatenated form ordered by activity ID must correspond exactly to the FM that is obtained when the entire log is considered and all direct successions are entered into a large matrix.
In our experimental evaluation in Section 5, we show (a) that both EdgeMiner and established central algorithms result in identical FMs and thereby process models and (b) that in four out of five datasets an intermediate FM already complies with the FM of the whole event log by more than 90% after less than 5% of the events of the datasets.
4.5. Distributed Mining with EdgeMiner
By distributing the computation of FMs, we can efficiently implement key process mining algorithms, such as, for example, the Alpha Miner, Heuristics Miner, and Inductive Miner, while also supporting online conformance checking in a scalable and privacy-preserving manner. Below, we outline how EdgeMiner distributes each of these algorithms.
4.5.1. Distributed Alpha Miner
The Alpha Miner relies on identifying direct successions in event logs (cf. Section 2). Hence, it builds on causal relations between activities. We leverage EdgeMiner to distribute the computation of the FM from which the Alpha Miner constructs the process model. The Alpha Miner expects an FM consisting only of ones and zeros, i.e., if there is directly-follows relation or not. As EdgeMiner counts the number of successions, we simply set FM entries greater than zero to one.
4.5.2. Distributed Heuristics Miner
In contrast to the Alpha Miner, the Heuristics Miner not only considers whether a direct succession is present, but also considers the frequency and dependency measure of activity relationships. With a request, EdgeMiner merges all partial FMs to a global FM at the requesting entity, which includes the frequencies of the direct successions. From these, it calculates the dependency measures of the activity relationships. Next, the Heuristics Miner constructs a process model based on these frequencies and dependency measures.
4.5.3. Distributed Inductive Miner
The Inductive Miner first constructs a directly-follows graph and then recursively splits it into subprocesses to derive a process tree as process model. To obtain the directly-follows graph the Inductive Miner needs the direct successions present in the event log as well as the set of start activities and the set of end activities. EdgeMiner offers the direct successions in the form of an FM along with the two sets. Just as for the Alpha Miner, the frequencies of the direct successions are not relevant, so that all entries greater than zero are set to one.
4.5.4. Distributed Footprint-Based Online Conformance Checking
Online conformance checking continuously monitors process executions and compares them against a reference model in real time. Using EdgeMiner, each node generates and maintains a local, partial FM that captures the current state of the process at that node. As new events are detected, the local FMs are updated, and each node can use its FM to perform conformance checking against a predefined process model.
5. Evaluation
@@
In this section, we evaluate EdgeMiner in terms of correctness, accuracy, and performance. We begin by discussing implementation, baselines and datasets in Section 5.1. Next, Section 5.2 analyzes these datasets. In Section 5.3, we experimentally evaluate the correctness of EdgeMiner. We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section 5.4. Section 5.5 estimates FM sizes during an algorithm run, and Section 5.6 reflects on the evaluation results.
5.1. Evaluation Setup
5.1.1. Implementation
@@
We implement EdgeMiner in Python and use a network of docker containers to simulate the distributed nature of EdgeMiner.
5.1.2. Baselines
@@
In our experiments, we compare EdgeMiner with and without MFP Requesting to two baselines: (1) ”Query All” represents the upper bound where all nodes are queried, while (2) a theoretical oracle-like approach serves as the lower bound. The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise. Further, we compare EdgeMiner to traditional process discovery algorithms (van der Aalst et al., 2004; Weijters et al., 2006; Leemans et al., 2013) to validate that the resulting FMs are identical.
5.1.3. Datasets
@@
We conduct experiments on the five rel-world datasets (event logs): Hospital Log (van Dongen, 2011), Sepsis Cases (Mannhardt, 2016) and an IoT Event Log for Process Mining in Smart Factories [^2] We use the MainProcess.xes file of the non-cleaned version. (Malburg et al., 2022). These process-mining event logs naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 (van Dongen, 2017) and the Road Traffic Fine Management Process (de Leoni and Mannhardt, 2015) datasets. In our setting, each activity in a dataset represents a node in an IoT network.
@@
Table 1. Dataset properties and EdgeMiner performance (Standard deviation in parentheses if not stated otherwise).
@@
5.2. Dataset Analysis
@@
We begin our experimental evaluation with an analysis of our real-world event logs, see Table 1. Our analysis shows that even in large datasets like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event. Further, our analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events. This characteristic benefits EdgeMiner, as self-loops do not require communication.
5.3. Experimental Correctness & Fitness
@@
In addition to the analytical correctness proof in Section 4.4, we experimentally validate that the resulting FMs of EdgeMiner are identical to the ones computed by today’s centralized algorithms333we compare implementations in pm4py (https://pm4py.fit.fraunhofer.de). We test both on randomly generated artificial datasets and common real-world event logs, see Section 5.1.3. For all datasets the resulting FMs are identical at the end of the log, see Fig. 7.
@@
Next, we examine how the merged FMs converge over time: Fig. 7 demonstrates that nearly all datasets reach 90% fitness within less than 5% of the trace Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, as it is quite small, see Table 1. The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events before requesting the first FM. Moreover, our analysis suggests that reducing the sliding window for discarding old data has minimal impact on the resulting FM.
5.4. Number of Queried Nodes
To assess EdgeMiner’s performance, we analyze its two phases separately. For the first experiments, we assume a batch size of one.
Figure 4. Average number of nodes queried with and without MFP Requesting, including standard error. MFP Requesting and knowledge of the start events reduce communication demands by a factor of 7.5 to 30 depending on the dataset.
[…]